### v4

# DecoderTracker: Decoder-Only Method for Multiple-Object Tracking

Abstract
Decoder-only models, such as GPT[1], have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer models. Over the years, end-to-end models based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However, the significant computational resource consumption of these models leads to less friendly inference speeds and training times. To address these issues, this paper attempts to construct a lightweight Decoder-only model: DecoderTracker for end-to-end multi-object tracking. Specifically, drawing on some real-time detection models, we have developed an image feature extraction network which can efficiently extract features from images to replace the encoder structure. In addition to minor innovations in the network, we analyze the potential reasons for the slow training of MOTR-like models and propose an effective training strategy to mitigate the issue of prolonged training times. On the DanceTrack dataset, without any bells and whistles, DecoderTracker’s tracking performance slightly surpasses that of MOTR, with approximately twice the inference speed. Furthermore, DecoderTracker requires significantly less training time compared to MOTR.
1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer[2] models in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT models like MOTR[3] have been introduced in the MOT domain. These models offer advantages over the widely used tracking-by-detection methods[4, 5, 6, 7], such as eliminating the need for manual design or selection of features. Additionally, some enhanced methods[8, 9, 10] of MOTR have demonstrated the potential of this model class in terms of tracking accuracy.
However, the inference speed of these models is often suboptimal, and their RNN[11] like linear inference structure makes them less GPU friendly during training.
[…]
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the model to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? Based on this idea, we proposed the Tracking Box Selection Process (TBSP).
Although the methods we propose have not yet enabled the inference speed of MOTR-like models to achieve real-time performance, we believe that this paper could provide valuable insights for future researchers.
In summary, the contributions of this paper are as follows:
  • 1)
    The primary contribution involves the development of a novel end-to-end tracking network, DecoderTracker, which is a decoder-only method. Compared to MOTR, this network achieves faster inference speeds and improved tracking performance.
  • 2)
    A second key innovation introduces a unique training strategy characterized by a three-stage process. This strategy accelerates model convergence, reducing overall training time and enhancing training process efficiency.
  • 3)
    The third contribution introduces TBSP, a strategy that can serve as a preliminary weakly supervised training approach for end-to-end MOT model training.
2 Related work
Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios. The track query set, initially empty, is dynamically adjusted in length. Detect queries play a role in identifying new objects.
2.1 MOTR Series
Unlike tracking-by-detection models exemplified by SORT[4, 14, 5], and previous transformer-based MOT models[15, 16], MOTR[3] is the first model to achieve end-to-end multi-object tracking. It extends the object query from DETR[17] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for modeling long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
MOTRv2[18] enhances the detection capabilities of the original MOTR by integrating the pre-trained object detector YOLOX[19]. This integration involves using YOLOX generated proposal boxes as anchors to initialize object queries, a method that not only bolsters detection accuracy but also helps to alleviate conflicts between detection and association tasks, all the while maintaining the query propagation features inherent to MOTR.
Similarly, MeMOT[10] innovates by incorporating track queries that model and predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
Moreover, MOTRv3[9] and CO-MOT[20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework. MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance. In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces "shadows" to enhance the diversity and quantity of queries, further refining the MOT capabilities.
2.2 Real Time Detection Method
When discussing real-time detection models, most people initially think of YOLO[21], which transformed the problem of object detection into a regression model, achieving efficient object detection. The model divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. The most advanced versions in the YOLO series currently are YOLOv8.2[22] and YOLOv9[23].
Recently, the real-time detection model based on DETR, known as RT-DETR[24], has also demonstrated its extraordinary performance. RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the model structure.
Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary. Therefore, inspired by the YOLO series[25, 22, 26], we designed a feature extraction network.
3 DecoderTracker
Figure 2: The overall structure of DecoderTracker. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. When the model is trained using the TALA strategy, the TBSP can be skipped.
Figure 3: The structure of IMFENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. For detailed comparisons, please refer to the appendix.
3.1 DecoderTracker Architecture
DecoderTracker’s overall architecture is illustrated in the provided Fig.2. In this setup, a video sequence is input to Image Feature Extraction Network(IMFENet) to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder. These queries interact with the image features in the decoder to generate hidden states for bounding box predictions. Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
The structure of IMFENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction. The reason for not adopting the latest YOLOv9 is due to its slower training speed compared to YOLOv8. Through IMFENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
3.2 Tracking Box Selection Process
Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2
The initial design of the TBSP was intended merely to equip a detection network with basic tracking functionality. We observed that when a detection network directly employs the MOTR[3] method to update track queries, the tracking boxes for the same object tend to increase over time. This accumulation was found to be due to each newly generated detect query invariably being assigned to an already tracked object. In the DETR, a detect query can be assigned to any object in the image, as label assignment is determined through bipartite matching between all detection queries and ground truth data. MOTR addresses this by utilizing tracklet-aware label assignment (TALA), which suppressively modulates the generation of detect queries similar to track queries during training, thus mitigating this issue.
Consequently, we designed a simple filtering mechanism, TBSP, which enables the detection network to possess some tracking capabilities without additional training. However, given the network’s inherently weak tracking ability and the lack of robustness in TBSP, the tracking performance at this stage is quite limited. Nevertheless, we also discovered that using TBSP as a preprocessing step for TALA, and training the network in this fashion, can lead to some unexpected benefits.
Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training. This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the model to optimize itself. For a discussion on why TBSP is effective, please refer to the appendix.
We know that both track query and detect query contain information about the object’s bounding box(BBox). Let’s denote the BBox in $\hat{\omega}_{tr}^{t-1}$ as $[BBox_{1}^{tr}$, $BBox_{2}^{tr},...,BBox_{n}^{tr}]$, where $n$ is the number of track queries at time $t-1$. Each query in $\hat{\omega}_{tr}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
For each element in $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$, calculate the Intersection over Union (IOU) with each element in $[BBox_{1}^{tr},BBox_{2}^{tr},...,BBox_{n}^{tr}]$. If the IOU is greater than $\delta_{iou}$, the BBox corresponding to that detect Query will be discarded. Finally, the remaining detect queries are denoted as $\omega_{\det_{f}}^{t-1}$.
Similarly, assuming the track query at frame $t$ (where $t>2$) is denoted as $\omega_{tr}^{t}$, it can be expressed as:
$\hat{\omega}_{tr}^{t}\,\,=\omega_{det_{f}}^{t-1}\cup\,\,\hat{\omega}_{tr}^{t-1}\,\,$ (1)
The feasibility of TBSP is illustrated in the Fig. 4. Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our model compared to MOTR, we have not adopted the labeling strategy used in CO-MOT[8] and MOTRv3[9].
3.3 Loss and Training Methods
Due to MOTR’s similarity to RNN in its inference approach, its batch size is limited to 1. This often results in suboptimal utilization of GPU resources during training. Additionally, MOTR learns both image and motion features simultaneously. To expedite model training, this paper adopts a three-stage training strategy. In the first stage, emphasis is placed on training the network to extract target image features efficiently. The second stage involves training the model on weak tracking aspects using TBSP as a pretraining step for the third stage. Lastly, the third stage employs the TALA strategy to comprehensively enhance the model’s tracking performance.
In the initial stage, the network is trained as a detection network, and the loss function is formulated defined as Eq.2:
$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)}{V_{i}}$ (2)
$\mathcal{L}$ is the loss of single frame. Notably, during this stage, we employed the IoU-aware Query Selection from RT-DETR, so the loss function utilized is from RT-DETR. Unlike the original DETR, this stage introduces a loss termed $\mathcal{L}_{enc}$. This loss involves treating the initialized queries as boxes for the purpose of loss computation. The specifics are outlined below:
$\mathcal{L}_{enc}(\hat{y},y)&=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},\hat{b},y,b)\\ &=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},c,IoU)\end{aligned}$ (3)
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively. Finally, the overall loss is as follows:
$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (4)
where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[27], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[28], $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients.
In this stage, the network is capable of initially learning the appearance features of objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Additionally, various data augmentation techniques [29, 30, 31] can be employed to expedite the model’s extraction of appearance features.
The second stage of training is conducted using MOTR’s Collective Average Loss. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as Eq.5.
$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\sum_{n=1}^{N}{\left(\mathcal{L}\left(\left.\widehat{Y}_{tr}^{t}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)+\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)\right)}}{\sum_{n=1}^{N}{\left(V_{t}\right)}},N=5$ (5)
where $V_{i}=V_{tr}^{i}+V_{det}^{i}$ denotes the total number of ground-truths objects at frame $i$. $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively.
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage.
The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
4 Experiments
4.1 Datasets and Metrics
To validate DecoderTracker’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[32], MOT17[33], and KITTI[34].
DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
MOT17 focuses on multiple object tracking in public spaces, primarily pedestrians. The dataset includes seven scenes, divided into training and testing clips, providing object detection annotations suitable for both online and offline tracking approaches.
KITTI serves as a research dataset for autonomous driving, encompassing stereo, optical flow, odometry, 3D object detection, and tracking tasks. Captured across diverse environments using advanced sensors, it provides raw data and task-specific benchmarks.
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[35, 36]. These metrics collectively provided a comprehensive and robust evaluation of our model’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our model’s tracking capabilities.
4.2 Implementation Details
The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline model surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [10], MOTRv2 [18], MOTRv3 [9], and CO-MOT [8].
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. One more thing, he image size for the input network is $640\times 640$.
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1. For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [37], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP. For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [38] were selected for additional training.
4.3 Performance Evaluation and Comparison
In Table 1, a comprehensive comparison of training time and speed was conducted between DecoderTracker and the MOTR series. It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations. This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
(a) Comparison of Training Time on MOT17
Method Device Training Time
MOTR[3] 8 2080ti $\sim$ 96.0 hours
DecoderTracker stage1 1 2080ti 12.1 hours
DecoderTracker stage2 1 2080ti 30.8 hours
DecoderTracker(TBSP) 1 2080ti 43.4 hours
DecoderTracker 1 2080ti 68.7 hours
(b) Comparison of Speed on Dancetrack
Method FPS
MOTR[3] 9.5
MOTRv2[18] 6.9
MOTRv3[9] 10.6
DecoderTracker(TBSP) 18.5
DecoderTracker 19.6
Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU. The data for the MOTR series is sourced from [3], [9] or their open source address.
DecoderTracker demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed. The primary reason behind this efficiency gain lies in the architectural simplicity of DecoderTracker. Unlike MOTR, DecoderTracker adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design. This reduction in complexity enables DecoderTracker to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU. In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training. The absence of an encoder in DecoderTracker contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[32] benchmark. Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR.
Methods HOTA IDF1 DetA AssA MOTA
Non-End-to-end:
ByteTrack[14] 47.3 52.5 71.6 31.3 89.5
GTR[39] 48.0 50.3 72.5 31.9 84.7
DST-track[40] 51.9 51.0 82.3 34.6 84.9
OC-SORT[41] 54.1 54.2 80.4 38.0 89.4
End-to-end:
MeMOTR[10] 63.4 65.5 77.0 52.3 85.4
MOTR[3] 54.2 51.5 73.5 40.2 87.6
DecoderTracker(Ours) 54.2 54.2 78.7 37.4 87.8
On the Dancetrack test benchmark, DecoderTracker(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3). Particularly noteworthy is the remarkable 19.6 FPS achieved by DecoderTracker, emphasizing its outstanding performance in inference speed. Table 2, Table 3, and Table 4 exhibit the final results of end-to-end DecoderTracker on the Dancetrack, MOT17, and KITTI benchmark, respectively. In contrast to MOTR, DecoderTracker demonstrates competitive performance across multiple metrics on both datasets. It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3. The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our DecoderTracker. MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training. Consequently, comparing its tracking performance directly with that of DecoderTracker and MOTR may not be entirely fair.
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17[33] set.
method HOTA IDF1 DetA AssA MOTA
Non-End-to-end:
FairMOT[42] 59.3 72.3 60.9 58.0 73.7
ByteTrack[14] 62.8 77.2 63.8 62.2 78.9
StrongSORT[39] 63.5 78.3 63.6 63.7 78.5
UTM[43] 64.0 78.7 65.9 62.5 81.8
OC-SORT[41] 63.2 77.5 63.2 63.2 78.0
Deep OC-SORT[6] 64.9 80.6 64.1 65.9 79.4
End-to-end:
MeMOTR[10] 58.8 69.0 59.6 58.4 72.5
MOTR[3] 57.2 59.4 58.9 55.8 71.9
DecoderTracker(Ours) 55.7 70.5 55.7 56.0 66.5
For Dancetrack (Table 2), DecoderTracker achieves comparable results to MOTR, with same scores in HOTA but outperforming MOTR in DetA. Specifically, DecoderTracker achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
In the MOT17 (Table 3), DecoderTracker demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA. Despite trailing slightly behind MOTR in these metrics, DecoderTracker compensates with a comparable DetA and excels in IDF1 by almost 10 points. This advantage can be attributed to DecoderTracker’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand. Additionally, DecoderTracker maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI[34] set.
Pedestrian Car
Tracker HOTA MOTA AssA HOTA MOTA AssA
Non-End-to-end:
CenterTr[44] 40.3 53.8 36.9 73.0 88.8 71.1
PermaTr[45] 47.4 65.0 43.6 77.4 90.8 77.6
PolarMOT[46] 43.6 46.9 48.1 75.1 80.0 76.9
TripletTrack[47] 50.0 46.5 57.8 73.6 84.3 74.7
End-to-end:
DecoderTracker 51.5 56.8 58.4 72.2 83.2 73.8
Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark. DecoderTracker achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position DecoderTracker as a strong contender in MOT on the KITTI.
4.4 Ablation Study
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR[48] or employing an MLP, on the model’s performance was investigated.
(a) Performance of DecoderTracker at different training stages
Training Stage HOTA AssA MOTA
First 38.5 21.6 76.8
Second 44.5 25.5 84.2
Final (TBSP) 51.9 34.6 84.9
Final 53.6 35.9 84.7
(b) The presence or absence of TAN during the second training stages can significantly impact the performance of the model.
Model HOTA AssA MOTA
Second 44.5 25.5 84.2
Second(w/o) 42.1 21.4 80.8
Final 53.6 35.9 84.7
Final(w/o) 53.4 36.1 84.5
(c) The impact of $\delta_{iou}$ on model performance when using TBSP
$\delta_{iou}$ HOTA AssA MOTA
0.3 49.7 32.3 79.1
0.5 51.9 34.6 84.9
0.7 51.3 34.8 83.2
0.9 50.0 32.8 81.3
(d) Effect of query position generation strategy
Setting HOTA MOTA FPS
sin/cos 53.3 85.1 18.5
MLP 53.6 84.7 19.6
Table 5: Ablation Studies on Our Proposed DecoderTracker on the DanceTrack validation set.
The performance of DecoderTracker improves consistently across different training stages, peaking in the final stage. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
It is noteworthy that in Table 5(b), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages. It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance. However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
5 Conclusion and Limitation
This paper presents an efficient decoder-only end-to-end MOT (Multi-Object Tracking) model that boasts faster inference speeds and reduced training times. Additionally, this paper introduces new training strategies aimed at accelerating model convergence. Although device limitations prevented us from validating our training strategies on the original MOTR model, we believe these strategies are effective. This research provides a more efficient benchmark model for the multi-object tracking field compared to MOTR, laying a foundation for the emergence of real-time end-to-end MOT models and offering insights for future development. In subsequent work, we should further explore more efficient network architectures to construct real-time MOT models. Of course, we also plan to incorporate techniques from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to enhance tracking performance.
Appendix A More Discusstion About TBSP
Figure 5: Refinement process for DecoderTracker on partially trained track queries. When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Fig. 1.
Table 6: Comparison of DecoderTracker with and without using TBSP on the Dancetrack val set.
Setting HOTA MOTA Epochs Total Time
TBSP 53.6 84.7 27 56.7 hours
w/o 53.4 84.2 40 85.2 hours
The initial reason for developing the MO-YOLO model was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
Fortunately, our DecoderTracker model has addressed this issue. Although its inference speed is not yet real-time, its convergence rate is already satisfactory. Although we have not achieved real-time inference speeds, we have significantly improved the speed of MOTR.
Regarding the TBSP, its original purpose was simply to endow the detection network with basic tracking functionality. However, it also provided some unexpected benefits, such as further reducing training time. Below, we list the results and required iterations if training was conducted directly using TALA, as shown in Table 6.
We believe the effectiveness of the TBSP strategy may be due to its function as a form of weak or self-supervised training for motion features, compared to TALA. Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence. For instance, in the NLP field, large language models like BERT[49] and GPT[1] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns. Subsequently, these models undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering. This approach has been empirically proven to significantly boost the performance of LLMs.
In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[50, 51]. Typically, these models are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
We believe that the TBSP, a weak supervision process, enables the model to preliminarily learn object motion information that would otherwise require many more training iterations. It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill. A similar situation may exist here.
Although TBSP did not enhance model performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
Appendix B Design of IMFENet
The design of this feature network is intended to enhance the model’s efficiency and increase its inference speed. Thus, our design primarily references the network structure of YOLOv8. The backbone structure of our model is almost identical to that of YOLOv8. Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M. This change is specifically reflected in the BottleNeckF module. In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load. These modifications are illustrated in Figure 6, and their impact on performance is shown in Table 7.
Figure 6: The difference between our proposed model and YOLOv8 in terms of modules. Our modules are more lightweight. Specific model details can refer to the figure.3.
Figure 7: DecoderTracker and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 benchmark.
Table 7: The impact of our network and YOLOv8 structure on DecoderTracker performance on the Dancetrack val set. Training time does not include the first stage
Setting HOTA IDF1 MOTA AssA FPS
IMFENet 53.6 53.8 84.7 35.9 19.6
YOLOv8 52.4 51.4 84.2 36.1 18.9
Appendix C Limitation
Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection models. Both models fail under similar circumstances, as illustrated in Figure.7. This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the model. Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR.