### v5

# DecoderTracker: Decoder-Only End-To-End method for Multiple-Object Tracking

Abstract
Decoder-only methods, such as GPT, have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer methods. Over the years, end-to-end methods based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However,The substantial computational resource consumption of these methods, coupled with the optimization challenges posed by dynamic data, results in less favorable inference speeds and training times. To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTrack, a novel end-to-end tracking method. Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTrack+ by incorporating a Fixed-Size Query Memory and refining certain attention layers. Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, with inference speeds 2.06 and 3.03 times faster than MOTR, respectively.
keywords:
Multi-object tracking , Decoder , End-to-end
journal: Pattern Recognition11footnotetext: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer [1] methods in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT methods like MOTR [2] have been introduced in the MOT domain. These methods offer advantages over the widely used tracking-by-detection methods [3, 4, 5, 6], such as eliminating the need for manual design or selection of features. Additionally, some enhanced methods [7, 8, 9] of MOTR have demonstrated the potential of this method class in terms of tracking accuracy.
[…]
Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios. The number of detect queries is typically fixed, while the number of track queries fluctuates throughout the tracking process. This discrepancy results in dynamic data during inference, which introduces optimization challenges that limit the actual runtime speed of the method. To address this, we propose the Fixed-Size Query Memory (FSQM), which pads the tracking query sequence with empty queries to maintain a consistent number of queries during inference. By eliminating dynamic data, we have developed DecoderTracker+, which has significantly improved inference speed compared to MOTR and DecoderTracker.
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the method to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? It is important to note that weak supervision, or self-supervised learning, has achieved significant success in many computer vision tasks [12, 13]. Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during the training. Thus, DecoderTracker has been constructed by us.
After completing the aforementioned architectural optimizations mentioned above, we observed that while the training duration was significantly optimized, the inference speed did not improve as expected. This phenomenon prompted us to delve deeper into the issue. In theory, streamlining the network architecture should more notably enhance inference speed rather than training efficiency. To investigate this anomaly, we systematically analyzed the time consumption distribution across various modules. The results revealed that the decoder module accounted for the majority of the time, consuming nearly two-thirds of the model’s total inference time, which is evidently unreasonable. Through further cross-model comparative analysis, we identified that this issue also exists in MOTR but is absent in its base model, Deformable DETR. Controlled variable experiments indicated that fixing the number of queries significantly reduces decoder latency. Considering factors such as GPU memory usage, we ultimately pinpointed the root cause to be the dynamic query mechanism. This mechanism introduces unnecessary overhead during inference, such as inefficient GPU memory allocation in modern deep learning frameworks ,e.g., PyTorch/TensorFlow, and may also hinder compilation optimizations like operator fusion due to dynamic computation graphs. To address this issue, we proposed the Fixed-Size Query Memory (FSQM) for handling fixed-query tracking. By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTrack+. Experimental results demonstrated that, although DecoderTrack+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
In summary, the contributions of this paper are as follows:
  • 1)
    The primary contribution involves the development of a novel end-to-end tracking network, DecoderTracker, which is a decoder-only method. Compared to MOTR, this network achieves faster inference speeds and improved tracking performance.
  • 2)
    The second key innovation involves the introduction of a unique training strategy characterized by a three-stage process. At the same time, we incorporated TBSP (Temporal Balanced Sampling and Propagation), a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods. These strategies accelerate method convergence, reduce overall training time, and enhance the efficiency of the training process.
  • 3)
    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTrack+, which addresses the additional latency caused by dynamic data. This provides a fully viable technical pathway for the engineering deployment of MOTR-like end-to-end MOT models.
2 Related work
2.1 MOTR Series
Unlike tracking-by-detection methods exemplified by SORT [3, 14, 4], and previous transformer-based MOT methods [15, 16], MOTR [2] is the first method to achieve end-to-end multi-object tracking. It extends the object query from DETR [17] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for methoding long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
MOTRv2 [18] enhances the detection capabilities of the original MOTR by integrating the pre-trained object detector YOLOX [19]. This integration involves using YOLOX generated proposal boxes as anchors to initialize object queries, a method that not only bolsters detection accuracy but also helps to alleviate conflicts between detection and association tasks, all the while maintaining the query propagation features inherent to MOTR.
Similarly, MeMOT [9] innovates by incorporating track queries that method and predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
MOTRv3 [8] and CO-MOT [20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework. MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance. In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces ”shadows” to enhance the diversity and quantity of queries, further refining the MOT capabilities.
2.2 Tracking by Detection
Tracking by Detection are among the most classical and widely used types of MOT. These methods typically employ a detector to first identify objects and then use various other methods for tracking.
In past years, the main methods within the Tracking by Detection category were SORT [3] and its derivative methods. The primary approach of these methods is to use a Kalman Filter to predict the positions of targets in the next frame and then perform data association with the detection results provided by the detector. Many subsequent improvements to these methods [5, 14, 4, 21] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association. There are also a series of methods like OC-SORT [22] that purely improve the Kalman Filter-based tracking method using kinematic methods.
In recent years, methods such as SUSHI [23], Glan [24] and PuTR [25] have emerged, which employ networks for tracking post-detection. The former two utilize Graph Neural Networks (GNN), while the last is based on transformer.
2.3 Real Time Detection Method
Figure 2: The overall structure of DecoderTracker. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. To put it simply, $q_{tr}$ represents the objects in the video sequence that need to be tracked continuously, while $q_{d}$ represents the new objects appearing in the video frame. When the method is trained using the TALA strategy, the TBSP can be skipped. The process of DecoderTracker+ is consistent with this scenario, but the main difference lies in the updating of $q_{tr}$ and whether it is of fixed length. The FSQM of DecoderTracker+ pads $q_{tr}$ with all-zero queries. This approach effectively addresses optimization issues caused by dynamic data.
When discussing real-time detection methods, most people initially think of YOLO [26], which transformed the problem of object detection into a regression method, achieving efficient object detection. The method divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. The most advanced versions in the YOLO series currently are YOLOv8.2 [27] and YOLOv9 [28].
Recently, the real-time detection method based on DETR, known as RT-DETR [29], has also demonstrated its extraordinary performance. RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the method structure.
Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary. Therefore, inspired by the YOLO series [28, 27], we designed a feature extraction network.
3 Method
3.1 DecoderTracker
Figure 3: The structure of FENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. Specific module details can be found in Figure 4. The extracted features will be fed into the Decoder and introduced to the cross-attention module to interact with the queries. Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder and the TAN.
3.1.1 DecoderTracker Architecture
DecoderTracker’s overall architecture is illustrated in the provided Figure 2. In this setup, a video sequence is input to image Feature Extraction Network(FENet) to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder. These queries interact with the image features in the decoder to generate hidden states for bounding box predictions. Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
3.1.2 Design of FENet
The structure of FENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction. The reason for not adopting the latest YOLOv9 or YOLOv10 is due to its slower training speed compared to YOLOv8. Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
The design of this FENet is intended to enhance the method’s efficiency and increase its inference speed. Thus, our design primarily references the network structure of YOLOv8. The backbone structure of our method is almost identical to that of YOLOv8. Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
Figure 4: Specific module details, modules not specified in the diagram are network structures or APIs that come with PyTorch. These modules are primarily constructed based on CNN, capable of efficiently extracting image features. The employment of residual connections ensures that issues like gradient explosion and vanishing gradients are largely avoided.
3.1.3 Tracking Box Selection Process
Algorithm 1 TBSP flow
0:  $\hat{\omega}_{tr}^{t-1}$: Track queries assignment, $\hat{\omega}_{det}^{t-1}$: Detect queries assignment, $\delta_{iou}$: IOU threshold
1:  $[BBox_{1}^{tr},BBox_{2}^{tr},\dots,BBox_{n}^{tr}]\leftarrow$ BBoxes from $\hat{\omega}_{tr}^{t-1}$
2:  $[BBox_{1}^{det},BBox_{2}^{det},\dots,BBox_{m}^{det}]\leftarrow$ BBoxes from $\hat{\omega}_{det}^{t-1}$
3:  for  $BBox_{j}^{det}$ in $[BBox_{1}^{det},\dots,BBox_{m}^{det}]$ do
4:     for  $BBox_{i}^{tr}$ in $[BBox_{1}^{tr},\dots,BBox_{n}^{tr}]$ do
5:        Compute $IOU(BBox_{j}^{det},BBox_{i}^{tr})$
6:        if $IOU>\delta_{iou}$ then
7:           Discard $BBox_{j}^{det}$
8:           break
9:        end if
10:     end for
11:  end for
12:  $\omega_{\det_{f}}^{t-1}\leftarrow$ Remaining detect queries assignment
The initial design of the TBSP was intended merely to equip a detection network with basic tracking functionality. We observed that when a detection network directly employs the MOTR [2] method to update track queries, the tracking boxes for the same object tend to increase over time. This accumulation was found to be due to each newly generated detect query invariably being assigned to an already tracked object. In the DETR, a detect query can be assigned to any object in the image, as label assignment is determined through bipartite matching between all detection queries and ground truth data. MOTR addresses this by utilizing tracklet-aware label assignment (TALA), which suppressively modulates the generation of detect queries similar to track queries during training, thus mitigating this issue.
Consequently, we designed a simple filtering mechanism, TBSP, which enables the detection network to possess some tracking capabilities without additional training. However, given the network’s inherently weak tracking ability and the lack of robustness in TBSP, the tracking performance at this stage is quite limited. Nevertheless, we also discovered that using TBSP as a preprocessing step for TALA, and training the network in this fashion, can lead to some unexpected benefits. Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training. This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the method to optimize itself.
Formally, we denote the predictions of track queries as $\widehat{Y}_{tr}$ and predictions of detect as $\widehat{Y}_{\text{det }}.Y_{\text{new }}$ is the ground-truths of newborn objects. The label assignment results for track queries and detect queries can be written as $\omega_{tr}$ and $\omega_{\text{det}}$. For frame $i$, label assignment for detect queries is obtained from bipartite matching among detect queries and newborn objects. We know that both track query and detect query contain information about the object’s bounding box (BBox). Let’s denote the BBox in $\hat{\omega}_{tr}^{t-1}$ as $[BBox_{1}^{tr}$, $BBox_{2}^{tr},...,BBox_{n}^{tr}]$, where $n$ is the number of track queries at time $t-1$. Each query assignment in $\hat{\omega}_{det}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
For each element in $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$, calculate the Intersection over Union (IOU) with each element in $[BBox_{1}^{tr},BBox_{2}^{tr},...,BBox_{n}^{tr}]$. If the IOU is greater than $\delta_{iou}$, the BBox corresponding to that detect Query will be discarded. Finally, the remaining detect queries are denoted as $\omega_{\det_{f}}^{t-1}$.
Similarly, assuming the track query at frame $t$ (where $t>2$) is denoted as $\omega_{tr}^{t}$, it can be expressed as:
$\hat{\omega}_{tr}^{t}\,\,=\omega_{det_{f}}^{t-1}\cup\,\,\hat{\omega}_{tr}^{t-1}\,\,$ (1)
The feasibility of TBSP is illustrated in the Algorithm. 1. Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our method compared to MOTR, we have not adopted the labeling strategy used in CO-MOT [7] and MOTRv3 [8].
Figure 5: Refinement process for DecoderTracker on partially trained track queries. When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Figure. 1. The process in DecoderTracker+ is similar, but there are some differences in the removal and addition of track queries. The detailed operations are discussed in the subsequent sections.
3.1.4 Query and Query Positions
MOTR was initially built upon Deformable DETR [11]. The original DETR was considered slow to converge due to its Hungarian matching algorithm. Although Deformable DETR significantly improved the convergence speed of DETR-like methods through architectural enhancements, subsequent works such as DN-DETR [30] and DAB-DETR [31] further corroborated the assertion that the Hungarian matching algorithm was indeed problematic. In DETR and Deformable DETR, the queries are typically initialized as a set of non-interpretable parameters, which is evidently suboptimal for training. Drawing inspiration from RT-DETR’s query generation approach, we directly generate queries from the input decoder features f. This process can be specifically represented as follows:
$q_{d}&=\text{LayerNorm}(\text{Linear}(f))\\ \end{aligned}$ (2)
One more thing, to expedite method convergence, we implement a Query Positions generation strategy akin to CO-MOT. Drawing inspiration from DAB-DETR [31], we employ a fixed transformation to convert reference points into Query Positions. While experimentation with the RT-DETR approach involving a Multilayer Perceptron (MLP) to transform reference points shows no significant performance discrepancy, it is important to note that the use of MLP substantially increases training time.
3.1.5 Loss and Training Methods
To expedite method training, this paper adopts a three-stage training strategy. In the first stage, emphasis is placed on training the network to extract target image features efficiently. The second stage involves training the method on weak tracking aspects using TBSP as a pretraining step for the third stage. Lastly, the third stage employs the TALA strategy to comprehensively enhance the method’s tracking performance.
In the initial stage, the network is trained as a detection network, and the loss function is formulated defined as Eq.3:
$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (3)
where $\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the frame predictions. The loss is calculated by ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$. $\mathcal{L}_{cls}$ denotes the focal loss[32], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[33]. $\mathcal{L}_{enc}$ is the anchor loss of DAB-DETR [31] and $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients, which are set to 1,2,5,5 in our method, just like DAB-DETR.
In this stage, the network is capable of initially learning the appearance features of objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Additionally, various data augmentation techniques [34, 35, 36] can be employed to expedite the method’s extraction of appearance features.
The second stage of training is conducted using MOTR’s Collective Average Loss. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as Eq.4.
[…]
where $V_{i}=V_{tr}^{i}+V_{det}^{i}$ denotes the total number of ground-truths objects at frame $i$. $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively. $\mathcal{L}$ is the loss of single frame, which is similar to Eq.3.
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage. Thanks to our various strategies and improvements to the network, we achieved huge improvements in training time and inference speed compared to MOTR.
The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the method, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
3.2 DecoderTracker+
Following the improvements in previous sections that led to the proposal of DecoderTracker, we observed suboptimal enhancements in inference speed, failing to meet anticipated performance benchmarks. Through latency profiling of individual modules, we identified the decoder as the primary bottleneck limiting acceleration. Subsequent in-depth analyses of GPU memory utilization and query quantity revealed that the core issue likely stems from dynamic data structures. Compared to static data configurations, dynamic data introduces inference challenges such as inefficient memory allocation scheduling, with empirical evidence provided in Figure 6.
To address this, we present DecoderTracker+ in this chapter. By implementing a Fixed-Size Query Memory(FSQM) mechanism that maintains a constant number of tracking queries, our approach enables deeper optimization compatibility with modern deep learning frameworks ,e.g., PyTorch, TensorFlow. Furthermore, the deterministic data structure facilitates straightforward conversion to TensorRT, paving the way for accelerated deployment of end-to-end MOT methods on edge devices.
DecoderTracker+ maintains the core architectural framework and operational workflow of its predecessor, as systematically illustrated in Figure 2, while introducing strategic innovations primarily in two key dimensions: Fixed-Size Query Memory and Attention Computation Optimization. Subsequent subsections provide detailed technical expositions of these advancements.
Figure 6: The two figures illustrate the changes in Decoder memory occupancy and the latency of various components for MOTR, its base model Deformable DETR, our DecoderTracker, and DecoderTracker+ on the MOT17-14 sequence. In the right figure, the backbone of our methods is represented by FENet, and since they do not have an Encoder, their latency on the Encoder is 0. The memory changes of DecoderTracker are basically consistent with those of MOTR, so they are not displayed here. It can be clearly seen from the figures that dynamic data leads to unstable memory occupancy of the Decoder and abnormally high latency.
3.2.1 Fixed-Size Query Memory
FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$, where $N$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension. Compared to dynamic query mechanisms, this approach introduces:
$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N}}_{\text{confidencevector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (5)
to model memory for the continuous tracking process. During initialization, set $q_{ftr}^{\mathcal{M}}=\mathbf{0}$ and $C_{\text{tr}}^{(0)}=\mathbf{0}$ for all queries, where all queries are initialized to zero to avoid interference with networks that do not rely on attention mechanisms. Additionally, all inactive queries are also initialized to zero. Moreover, a global ID pool, denoted as $\mathcal{ID}$, is introduced. Each active tracking query is assigned a unique ID that accompanies the query as it moves across video frames. When a query is no longer associated with a tracked object, its corresponding ID is deleted and will not be reused. A new ID is assigned to new active queries. By the way, all inactive queries are assigned an ID of -1 and their bounding boxes are initialized with zero values.
During each frame processing, iterative updates are performed as follows:
  1. 1.
    New query injection: For detection queries assignment $\{\hat{\omega}_{\text{det}}^{(i)}\}$ with detection confidence $c_{i}>\tau_{\text{in}}$, each query is directly assigned the first available ID from the global ID pool (which initially contains the value -1 for new queries). The corresponding query in the memory pool is updated using a momentum update rule:
    $j=\mathop{\arg\min}\limits_{k}(\mathcal{ID}[k]=-1),$ (6)
    $\hat{\omega}_{tr}^{\mathcal{M}}{(t)}[j]=\hat{\omega}_{\text{det}}^{(i)}$
    After updating the query, its corresponding ID is assigned from the global pool.
  2. 2.
    Query removal: For the existing queries in the memory pool, confidence is monitored over time. If a query’s confidence falls below a threshold $\tau_{\text{out}}$ for three consecutive frames, it is considered for removal. Upon removal, the query is reset to zero, and its corresponding ID is set to -1. Specifically:
    $\text{If }C_{\text{tr}}^{(t-3)}[n]<\tau_{\text{out}},C_{\text{tr}}^{(t-2)}[n]<\tau_{\text{out}},C_{\text{tr}}^{(t-1)}[n]<\tau_{\text{out}},$ (7)
    then the query is removed by setting:
    $q_{tr}^{\mathcal{M}}[n]=\mathbf{0},\quad\mathcal{ID}[n]=-1.$ (8)
    where $S(\cdot)$ represents the prediction score function. The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
It is also important to emphasize that, aside from this, the processing of queries in FSQM is consistent with that in DecoderTracker. The specific process will not be repeated here, but can be referred to in the previous sections.
Algorithm 2 FSQM Online Update Procedure
0:  New query set $\hat{\omega}_{\text{det}}=\{q_{i},c_{i},BBox_{i}\}_{i=1}^{M}$, memory pool $\hat{\omega}_{tr}^{\mathcal{M}}=(q_{\text{tr}},C_{\text{tr}},\mathcal{ID},BBox_{tr})$
0:  Updated memory pool $\mathcal{M}^{\prime}$
1:  $\hat{\omega}_{\text{valid}}\leftarrow\{\hat{\omega}_{i}|c_{i}>\tau_{\text{in}}\}$Filter valid new queries}
2:  if $\hat{\omega}_{\text{valid}}\neq\emptyset$ then
3:     for $q_{j}\in\mathcal{\hat{\omega}}_{\text{valid}}$ do
4:        Assign the first available ID from the global ID pool $\mathcal{ID}$ to $q_{j}$
5:        $k\leftarrow\mathop{\arg\min}\limits_{n}\mathcal{ID}[n]=-1$Find the first inactive query}
6:        $q_{\text{tr}}[k]\leftarrow q_{j},BBox_{\text{tr}}[k]\leftarrow BBox_{j}$Momentum update}
7:        $C_{\text{tr}}[k]\leftarrow c_{i}$Reset confidence to 1.0}
8:     end for
9:  end if
10:  for $n\in\{1,...,N\}$ do
11:     if $C_{\text{tr}}[n]<\tau_{\text{out}}$ for 3 consecutive frames then
12:        $q_{\text{tr}}[n]\leftarrow\mathbf{0},BBox_{\text{tr}}[k]\leftarrow\mathbf{0}$Set query and BBox to zero (inactive)}
13:        $\mathcal{ID}[n]\leftarrow-1$Reset ID to -1}
14:     end if
15:  end for
16:  return $\mathcal{M}^{\prime}=(Q_{\text{tr}},C_{\text{tr}},\mathcal{ID})$
3.2.2 Attention Mechanism Optimization
Since some queries in FSQM are empty, we need to explore the network structure to ensure that these empty queries do not affect the final results. The network processing these queries consists only of linear layers, normalization layers, dropout, activation function, cross-attention with image features, and finally, self-attention. Our first step is to analyze how empty queries may impact the output of these networks.
Firstly, we will analyze some unaffected networks. These network architectures inherently avoid commingling active and inactive queries in their computational processes. Therefore, during inference, there is no risk of interference arising from their interactions. Our subsequent analysis need only focus on verifying whether these networks produce gradients during the training.
For linear layers, zero-valued queries produce the following output:
$W\mathbf{0}+b=b$ (9)
Here, $W$ is the weight matrix, $\mathbf{0}$ is the zero-valued query, and $b$ is the bias term. Since the zero-valued query contributes nothing, only the bias remains, which can be neutralized through layer normalization. After applying layer normalization, the output is given by:
$\hat{y}=\frac{y-\mu}{\sigma}\cdot\gamma+\beta$ (10)
where $y=b$, $\mu$ and $\sigma$ are the mean and standard deviation of the output across the batch, and $\gamma$ and $\beta$ are learnable parameters. Since the bias term $b$ is small in practice, empirical tests show that its $L_{2}$ norm $\|b\|_{2}$ is very small, with $\|b\|_{2}<0.03$ in our implementation.
For normalization layers, since the zero-valued queries do not produce any meaningful activation, they do not affect the overall normalization process. The query’s zero value is normalized along with other queries, resulting in minimal influence on the normalized output. This can be expressed as:
$q_{i}=\frac{q_{i}-\mu}{\sigma}\cdot\gamma+\beta\quad\text{for active queries }q_{i}$ (11)
where $\hat{q}_{i}$ is the normalized output for an active query $q_{i}$. For inactive queries (those initialized to zero), the effect is negligible due to their zero contribution.
For dropout, which randomly sets a fraction of the inputs to zero during training, inactive queries do not contribute to the dropout process. Inactive queries, having values of zero, remain unaffected by dropout, and the overall output for those queries is still zero. Mathematically, dropout can be expressed as:
$\text{Dropout}(\omega_{i})=\begin{cases}0&\text{if query is inactive (}q_{i}=0\text{)}\\ q_{i}\cdot d&\text{if query is active}\end{cases}$ (12)
where $d\in[0,1]$ is a random variable representing the dropout rate. Inactive queries (zero-valued) remain zero regardless of whether dropout is applied.
For activation function, since only GELU is used to process queries, we will only discuss GELU here. In the case of GELU, the output for a zero-valued query is:
$\text{GELU}(q_{i})=0.5\omega_{i}\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}(q_{i}+0.044715q_{i}^{3})\right)\right)\quad\text{for }q_{i}=0\implies\text{GELU}(0)=0.$ (13)
Thus, for activation function, zero-valued queries result in zero output, ensuring that they do not influence the subsequent layers.
For cross-attention, given key matrix $K$ and dimension $d$ of Query, the attention scores become:
$\text{Attn}(Q_{\text{zero}},K)=\frac{\mathbf{0}\cdot K^{T}}{\sqrt{d}}=\mathbf{0}$ (14)
resulting in zero contribution after softmax normalization.
In summary, for linear, normalization, dropout, activation layers and cross-attention, zero-valued queries contribute nothing to the computation, ensuring that they do not interfere with the active queries during the forward and zero-valued in the forward inherently prevent gradient computation, thereby eliminating any impact on the training.
Regarding the self-attention layers, since the queries, keys, and values used in computing self-attention are all derived from the queries themselves, inactive queries inevitably introduce interference to the output. While a detailed analysis of this phenomenon is omitted here, we address this issue effectively by implementing a masking mechanism. Here we use the self attention layer in Decoder to illustrate, and the similarity in TAN will not be further elaborated.
When calculating self-attention, we first compute an attention weight matrix $A$:
$A=\text{softmax}\left(\frac{q(q^{\prime})^{T}}{\sqrt{d}}+\log(\text{Mask})\right)$ (15)
where $q\in\mathbb{R}^{2N\times d}$ is obtained by concatenating $q_{tr}^{\mathcal{M}}$ from FSQM and $q_{det}$, i.e., $q=\text{concat}(q_{tr}^{\mathcal{M}},q{det})$ in first decoder layer, while in other layers, $q$ is the output from the previous layer. $(\text{Mask}\in\mathbb{R}^{2N\times 2N}$ is the mask we need to generate. Its dimensions are the same as those of $A$, and it ensures that after applying the softmax function, the unwanted values are set to zero. By appropriately using this mask during the final self-attention output $A\times V$ where $V$ is derived from $Q$ with positional encoding, the effect of empty queries on the final result can be effectively eliminated.
Here, we extend $\mathcal{ID}$ and then generate the mask.
$\mathcal{ID}_{\text{ext}}=\text{Concat}(\mathcal{ID}_{\text{tr}},\mathbf{0}_{N})$ (16)
where, $\mathbf{0}_{N}$represents a fixed identifier for detect queries.
$\text{Mask}_{ij}=\begin{cases}0&\text{if }\mathcal{ID}_{ext}[i]=-1\lor\mathcal{ID}_{ext}[j]=-1\\ 1&\text{otherwise}\end{cases}$ (17)
By implementing this mask mechanism, we ensure that values associated with inactive queries in matrix $A$ are nullified to zero, thereby eliminating their influence on the final output. Notably, as demonstrated through our preceding analyses, the inactive query positions inherently maintain zero values across all decoder layers. Consequently, this masking mechanism uniformly applies to self-attention operations in every decoder layer without requiring layer-specific adaptations.
In terms of the mask in the self-attention layer of the TAN, the generation process and mechanism are largely consistent with those described earlier; however, the dimensions in TAN are fewer, and there is no need to expand $\mathcal{ID}$.
4 Experiments
4.1 Datasets and Metrics
To validate DecoderTracker and DecoderTracker+’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[37], MOT17[38], and KITTI[39]. DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
MOT17 focuses on multiple object tracking in public spaces, primarily pedestrians. The dataset includes seven scenes, divided into training and testing clips, providing object detection annotations suitable for both online and offline tracking approaches.
KITTI serves as a research dataset for autonomous driving, encompassing stereo, optical flow, odometry, 3D object detection, and tracking tasks. Captured across diverse environments using advanced sensors, it provides raw data and task-specific benchmarks.
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[40, 41]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our method’s tracking capabilities.
4.2 Implementation Details
The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline method surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [9], MOTRv2 [18], MOTRv3 [8], and CO-MOT [7].
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. One more thing, the image size for the input network is $640\times 640$.
(a) Comparison of Training Time on MOT17
Method Device Training Time
MOTR[2] 8 2080ti $\sim$ 96.0 hours
DecoderTracker stage1 1 2080ti 12.1 hours
DecoderTracker stage2 1 2080ti 30.8 hours
DecoderTracker(TBSP) 1 2080ti 43.4 hours
DecoderTracker 1 2080ti 68.7 hours
(b) Comparison of Speed on Dancetrack
Method FPS
MOTR[2] 9.5
MOTRv2[18] 6.9
MOTRv3[8] 10.6
DecoderTracker(TBSP) 18.5
DecoderTracker 19.6
DecoderTracker+ 28.8
Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU. The training time of DecoderTracker+ is basically the same as that of DecoderTracker, so it will not be listed separately here.
Figure 7: IDF1-MOTA-HOTA and IDF1-MOTA-FPS comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on Dancetrack test sets. For the left image, the horizontal axis is MOTA, the vertical axis is IDF1, and the radius of the circle is HOTA. For the right image, the horizontal axis is FPS, the vertical axis is IDF1, and the radius of the circle is HOTA. ”*” represents additional data was used for training. Our DecoderTracker+* achieves the best IDF1 and HOTA and comparable speed performance.
Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[37] test set. Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR. * means that extra data was used during training.
Methods IDF1 HOTA AssA DetA MOTA
Non-End-to-end:
TransTrack[15] 45.2 45.5 27.5 75.9 87.7
GTR[21] 50.3 48.0 31.9 72.5 84.7
DST-track[42] 51.0 51.9 34.6 82.3 84.9
ByteTrack[14] 52.5 47.3 31.3 71.6 89.5
OC-SORT[22] 54.2 54.6 38.0 80.4 89.4
Deep OC-SORT[5] 61.5 61.3 45.8 82.2 92.3
End-to-end:
MOTR[2] 51.5 54.2 40.2 73.5 79.7
DecoderTracker 54.2 54.2 35.7 78.7 87.8
DecoderTracker+ 54.4 54.5 35.6 79.0 88.1
DecoderTracker* 62.9 60.8 47.1 81.2 91.1
DecoderTracker+* 62.2 61.2 46.0 81.8 91.9
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [38] set. * means that extra data was used during training.
method IDF1 HOTA DetA AssA MOTA
Non-End-to-end:
YOLOTracker[43] 65.1 - - - 71.9
FairMOT[44] 72.3 59.3 60.9 58.0 73.7
MSPNet[45] 74.6 - - - 71.9
ByteTrack[14] 77.2 62.8 63.8 62.2 78.9
OC-SORT[22] 77.5 63.2 63.2 63.2 78.0
StrongSORT[21] 78.3 63.5 63.6 63.7 78.5
UTM[46] 78.7 64.0 65.9 62.5 81.8
Deep OC-SORT[5] 80.6 64.9 64.1 65.9 79.4
End-to-end:
MOTR[2] 59.4 57.8 60.3 55.7 73.4
DecoderTracker 70.5 55.7 55.7 56.0 66.5
DecoderTracker+ 70.4 55.8 55.2 55.8 66.7
DecoderTracker* 77.2 61.8 62.1 60.1 76.8
DecoderTracker+* 77.5 62.2 62.5 60.7 76.7
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1.
For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [47], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. And, to address the issue of the limited scale of the MOT17 dataset, this paper employs a hybrid dataset. It integrates the MOT20 dataset and the complete CrowdHuman dataset, which has been spatially shifted, to expand data diversity. The amount of new data significantly exceeds the benchmark set (by approximately 4.3 times). To balance training bias, a stratified sampling mechanism was designed: within each epoch, samples from the MOT17 dataset are resampled twice, while samples from the supplementary sets are sampled once. This approach not only alleviates data skewness but also enhances the learning of benchmark set features, ensuring that the model maintains its original performance advantages while improving its adaptability to dense scenes.
For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP. To better compare with the state-of-the-art (SOTA) methods, when testing on the test set, we also used the CrowdHuman dataset as supplementary data and the validation set of Dancetrack was also used.
For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [48] were selected for additional training.
4.3 Main Result
The comprehensive benchmarking in Table 1 demonstrates DecoderTracker’s dual superiority in training efficiency and inference speed over MOTR-series methods. This design achieves 68.7 training hours on a single GPU, equivalent to 7.1× higher GPU-hour efficiency versus MOTR’s 768 GPU-hour consumption (8 GPUs × 96 hours). The phased training regimen—12.1 hours for feature embedding, 30.8 hours for temporal modeling, and 43.4 hours for TBSP-enhanced prediction—systematically constructs tracking capabilities while maintaining 83% shorter wall-clock time than competitors. Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS,
For DanceTrack ,the experimental results shown in Table 2 reveal that DecoderTracker and its enhanced variant DecoderTracker+ achieve comparable precision levels across core evaluation metrics. Both models demonstrate superior detection accuracy over existing end-to-end methods, with DetA scores of 78.7 and 79.0 respectively, significantly outperforming MOTR’s 73.5. Their MOTA metrics (87.8 and 88.1) approach SOTA non-end-to-end methods like ByteTrack (89.5), while maintaining balanced HOTA performance at 54.2 and 54.5. This parity suggests that architectural refinements between the two variants primarily yield marginal improvements rather than fundamental performance shifts. Notably, both configurations exhibit identical IDF1 scores (54.2 and 54.4), confirming consistent identity preservation capabilities. The minimal performance delta between variants implies optimized parameter utilization rather than architectural overhauls. When extended with supplementary training data (DecoderTracker* variants), both models achieve substantial gains, attaining 91.1-91.9 MOTA and 60.8-61.2 HOTA. While their AssA scores (35.6-35.7) remain below MOTR’s 40.2, the detection-centric advantages manifest through DetA superiority over all end-to-end and most non-end-to-end competitors. These findings position the DecoderTracker framework as a robust baseline for end-to-end tracking, achieving detection reliability comparable to sophisticated multi-stage systems while preserving model efficiency.
In the MOT17 benchmark (Table 3), DecoderTracker shows strong performance, especially in complex tracking scenarios. While slightly behind MOTR in HOTA and AssA, it compensates with comparable DetA and excels in IDF1 by nearly 10 points. This advantage is due to its simplified network structure, which achieves competitive accuracy with less training data. Moreover, DecoderTracker maintains robust overall performance, as shown by its high MOTA score, making it suitable for real-world applications. When additional training data is introduced, DecoderTracker’s performance further improves. For instance, DecoderTracker* achieves a significant boost in IDF1 to 77.2, HOTA to 61.8, and MOTA to 76.8. Similarly, DecoderTracker+* also shows substantial gains. This clearly demonstrates the positive impact of more training data on DecoderTracker’s tracking capabilities, highlighting its potential for even better performance with increased data support.
Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI [39] set.
Pedestrian Car
Tracker HOTA MOTA AssA HOTA MOTA AssA
Non-End-to-end:
CenterTr[49] 40.3 53.8 36.9 73.0 88.8 71.1
PermaTr[50] 47.4 65.0 43.6 77.4 90.8 77.6
PolarMOT[51] 43.6 46.9 48.1 75.1 80.0 76.9
TripletTrack[52] 50.0 46.5 57.8 73.6 84.3 74.7
End-to-end:
DecoderTracker 51.5 56.8 58.4 72.2 83.2 73.8
DecoderTracker+ 51.1 56.4 58.8 72.6 82.1 73.9
Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark. DecoderTracker achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position DecoderTracker as a strong contender in MOT on the KITTI.
Figure 8: The visualisation results of our proposed DecoderTracker on various datasets show that DecoderTracker performs well both under long time series and short time series, and also has good tracking performance for different kinds of objects. The visualization results of DecoderTracker and DecoderTracker+ are essentially identical, so we won’t display the results of D+ here.
4.4 Ablation Study Of DecoderTracker
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR [31] or employing an MLP, on the method’s performance was investigated.
(a) Performance of DecoderTracker at different training stages
Training Stage IDF1 HOTA MOTA
First 38.6 38.5 76.8
Second 44.3 44.5 84.2
Final (TBSP) 51.7 51.9 84.9
Final 53.4 53.6 84.7
(b) The impact of $\delta_{iou}$ on method performance when using TBSP
$\delta_{iou}$ IDF1 HOTA MOTA
0.3 50.1 49.7 79.1
0.5 51.7 51.9 84.9
0.7 51.3 51.3 83.2
0.9 49.8 50.0 81.3
(c) The presence or absence of TAN during the second training stages can significantly impact the performance of the method.
method IDF1 HOTA MOTA
Second 44.4 44.5 84.2
Second(w/o) 41.9 42.1 80.8
Final 53.4 53.6 84.7
Final(w/o) 53.2 53.4 84.5
(d) Effect of query position generation strategy
Setting IDF1 HOTA FPS
sin/cos 85.1 53.3 18.5
MLP 84.7 53.6 19.6
Table 5: Ablation Studies on Our Proposed DecoderTracker on the DanceTrack validation set.
The performance of DecoderTracker improves consistently across different training stages, peaking in the final stage. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and method configurations, showcasing its adaptability and versatility in MOT field.
It is noteworthy that in Table 5(c), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages. It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance. However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
4.4.1 More Discusstion About TBSP
The initial reason for developing the DecoderTrack method was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
Fortunately, our DecoderTracker method has addressed this issue. Although its inference speed is not yet real-time, its convergence rate is already satisfactory. Although we have not achieved real-time inference speeds, we have significantly improved the speed of MOTR.
Regarding the TBSP, its original purpose was simply to endow the detection network with basic tracking functionality. However, it also provided some unexpected benefits, such as further reducing training time. Below, we list the results and required iterations if training was conducted directly using TALA, as shown in Table 6.
We believe the effectiveness of the TBSP strategy may be due to its function as a form of weak or self-supervised training for motion features, compared to TALA. Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence. For instance, in the NLP field, large language methods like BERT[53] and GPT[54] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns. Subsequently, these methods undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering. This approach has been empirically proven to significantly boost the performance of LLMs.
Table 6: Comparison of DecoderTracker with and without using TBSP on the Dancetrack val set.
Setting IDF1 HOTA Epochs Total Training Time
TBSP 54.2 53.6 27 56.7 hours
w/o 53.5 53.4 40 85.2 hours
In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[12, 13]. Typically, these methods are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
We believe that the TBSP, a weak supervision process, enables the method to preliminarily learn object motion information that would otherwise require many more training iterations. It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill. A similar situation may exist here.
Although TBSP did not enhance method performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
4.5 Ablation Study Of DecoderTracker+
As can be seen from the results in Tables 2, 3, and 4, the performance of DecoderTracker and DecoderTracker+ is almost the same. This is not difficult to understand. In Section 3.3, we discussed that empty queries do not affect the network’s output results after reasonable mask generation. DecoderTracker+ does not improve the network; instead, it addresses the optimization issues caused by dynamic data. At the same time, to better illustrate the effectiveness of the generated masks, we provided the performance of DecoderTracker+ with and without masks on the Dancetrack val set in Table 7. The results show that when masks are not used, the inactive queries still have an impact on the final tracking results. Combining this with previous data, both theoretical and practical evidence demonstrates that the use of masks does not affect tracking performance when FSQM is increased after reasonable mask generation.
Table 7: The impact of mask on the performance of DecoderTracker+ on the Dancetrack val set.
mask IDF1 HOTA MOTA AssA
53.6 53.4 84.9 36.0
w/o 46.2 45.8 78.3 31.7
4.5.1 Experimental Analysis of FENet
Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M. This change is specifically reflected in the BottleNeckF module. In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load. These modifications are illustrated in Figure 9, and their impact on performance is shown in Table 8.
Figure 9: The difference between our proposed method and YOLOv8 in terms of modules. Our modules are more lightweight. Specific method details can refer to the Figure.3. Additionally, it should be noted that there is not much difference between this part of YOLO11 and YOLOv8, so no comparison has been made
Table 8: The impact of our network and YOLOv8 structure on DecoderTracker and DecoderTracker+ performance on the Dancetrack val set.
Setting IDF1 HOTA MOTA AssA FPS
DecoderTracker:
FENet 53.8 53.6 84.7 35.9 19.6
YOLOv8 51.4 52.4 84.2 36.1 18.9
1.88 DecoderTracker+:
FENet 53.6 53.4 84.9 35.7 28.2
YOLOv8 52.1 53.2 85.4 36.8 26.7
5 Conclusion and Limitation
Figure 10: DecoderTracker and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 benchmark.
This paper proposes an efficient decoder-only end-to-end MOT method, which achieves faster inference speed and shorter training time compared to the baseline MOTR. Additionally, we introduce new training strategies aimed at accelerating the convergence of the method. Due to hardware limitations, we were unable to verify our training strategies on MOTR, but we believe these strategies are generalizable. Finally, we conducted a detailed and in-depth analysis of the slow inference process in MOTR, finding that dynamic data significantly impacts inference speed. Subsequently, we applied FSQM to address this issue, greatly enhancing the inference speed of the tracking-by-query end-to-end paradigm. Overall, this study provides a more efficient benchmark method for the multi-object tracking field, lays the foundation for the emergence of real-time end-to-end MOT methods, and offers insights for future development. In future work, we should further explore more efficient network architectures to build real-time MOT methods. Of course, we also plan to integrate the technologies from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to improve tracking performance.
Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection methods. Both methods fail under similar circumstances, as illustrated in Figure.10. This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the method. Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR. Although we attempted to use a mixed dataset to address this issue, our method still underperforms compared to SOTA tracking-by-detection methods. How to resolve this issue will also be a key focus of future research.